OpenStudy (braydonlevi99):

Simplify The Radical:

OpenStudy (braydonlevi99):

\[5\sqrt{20} + \sqrt{24} - \sqrt{180} + 7\sqrt{54}\]

OpenStudy (_citrus_):

You're just simplifying the expression right? So it would be 4\[4\sqrt{5} + 23\sqrt{6}\]

OpenStudy (braydonlevi99):

How would you simplify them though?

OpenStudy (_citrus_):

ignore that one 4 next to be

TheSmartOne (thesmartone):

You need to simplify all the radicals and then add/subtract like terms. Example: \(\Large\sqrt{90} = \sqrt{9\times 10} = \sqrt{9}\cdot\sqrt{10} = 3\sqrt{10}\)

jhonyy9 (jhonyy9):

Levi do you understand what Thesmartone wrote above like an example ? so in this way sqrt20 = sqrt(5*4)=2sqrt5 simplified and so 5sqrt20 = 5*2sqrt5 = 10sqrt5 try the next in this way